OpenStudy (anonymous):

use distributive property to simplify 6(23)

OpenStudy (anonymous):

6 x 23 in simpler terms.

OpenStudy (asnaseer):

you should make the multiplication simpler by distributing it as follows: 6(23) = 6(20 + 3) = 6*20 + 6*3 = 120 + 18 = 138

OpenStudy (anonymous):

Dang, nice trick. Never knew about that.